Presentation with dviout

Toshio OSHIMA January 1, 2000

§0 Presentation by TEX

Many technical documents including mathematical formulas are written by TEX and they can be used for presentations by dviout for Windows. Hit [Space] key! (⇒ next page) Study the solution u(x) of the Shrödinger equation

(−\frac{1}{2} _{1≤j≤n}^{∑} \frac{∂^{2}}{∂x_{j}^{2}} + U(x)) u(x) = λu(x)

with the potential function

U(x) = _{1≤i<j≤n}^{∑} \frac{C}{(x_{i}−x_{j})^{2}.} \frac{C}{sinh^{2}(x_{i} − x_{j})} .

It can be written by Gauss hypergeometric function when n = 2. Bessel function when n = 2.

One may try to present this document by a key or mouse operation. If it is displayed under a cover sheet,

• push [Space] key to proceed step by step,

• move the mouse under pushing its left button.

For this purpose, open slisamp2.dvi (a sample with the slightly yellow background).

Or open slisamp3.dvi (without the cover sheet).

Otherwise jump to sample.dvi.

Note that the [ESC] key is the toggle switch between the presentation mode and the normal mode.